If the government saw that consumer confidence was high, what step can it take to shift the AD to the left?

a. Congress can pass tax cuts.
b. Government can increase its spending.
c. The Federal Reserve can increase interest rates.
Answer #1

Answer

The correct answer is (c) The Federal Reserve can increase interest rates

In order to shift AD to the left, Government has to use such policy at which for any given price Aggregate Expenditure will decrease.

Suppose Congress cuts taxes then. Note: AE(Aggregate planned expenditure) = C + I + G + NX

C = Co + c(Y - T) where C = Consumption , T = Tax , Co = Autonomous consumption, we can say that as Tax (T) decreases C increases => Aggregate expenditure will increase => AD will shift to the Right.

If Government increase its spending then G = government Spending will increase => AE = C + I + G + NX will increase => AD will shift to the right.

Now If The Federal Reserve increases interest rates. As interest rate increases investment decreases => I(Investment) will decrease =>  AE = C + I + G + NX will decrease => AD will shift to the Left.

Hence, the correct answer is (c) The Federal Reserve can increase interest rates